OpenStudy (anonymous):

Find the differential equation corresponding to the family of curves 4y=c(c-2x)^2 where c is an arbitrary constant.

OpenStudy (anonymous):

4(dy/dx) = -4c(c-2x)

OpenStudy (anonymous):

i took the derivative with respect to x

OpenStudy (anonymous):

\[\left(\begin{matrix}dy \\ dx\end{matrix}\right)^3 = 8xy \frac{ dy }{ dx}-16y^2\] this is answer

OpenStudy (anonymous):

solce this in details

OpenStudy (calculusfunctions):

\[4y =c(c -2x)^{2}\] \[4\frac{ dy }{ dx }=2c(c -2x)(-2)\] \[4\frac{ dy }{ dx }=-4c(c -2x)\] \[\frac{ dy }{ dx }=-c(c -2x)\] Sorry, It took long because I was busy doing other things while typing the solution, but it seems as though you know how to finish it.

OpenStudy (calculusfunctions):

You do know how to finish it right?

OpenStudy (anonymous):

no i don't understand please give full solution in detail

OpenStudy (calculusfunctions):

OK! I just have to take care of something personal first and then I'll type the solution when I return. I'll be back in a few minutes. I'll remain logged on.

OpenStudy (calculusfunctions):

I'm back. Are you sure you've written the question correctly? Is there any other given information? And the answer you gave, is the actual answer?

OpenStudy (anonymous):

ya i'm sure that question is correct and that answer is actual answer

OpenStudy (anonymous):

he gave the same answer that i gave